OpenStudy (anonymous):

P=12x+20y in. find an expression for the area of a square with the given perimeter?

OpenStudy (anonymous):

Perimeter of square = 4*Side 4S = 12x+20y S = 3x+5y Now, the area of a square is Side * Side Area = S*S =(3x+5y)^2

OpenStudy (anonymous):

ok thankks you